OpenStudy (anonymous):

what is the sum of the roots of the polynomial shown below? f(x)=3x^3+12x^2+3x-18

OpenStudy (campbell_st):

well given the polynomial \[P(x) =ax^3 + bx^2 + cx + d\] and if it has roots \[\alpha, \beta, \gamma \] then the sum of the roots is \[\alpha + \beta+ \gamma = \frac{-b}{a}\] in your question a = 3 and b = 12 hope it helps

OpenStudy (paxpolaris):

\[f(x)=3\left( x^3+4x^2+x-6 \right)\] f(1)=0 

OpenStudy (paxpolaris):

\[f(x)=\left( x-1 \right)\left( x^2+5x+6 \right)\\ f(x)=(x-1)(x+2)(x+3)\] 1-2-3 = -4 .. same as campbell_st ... answer verified